OpenStudy (iwanttogotostanford):

I KEEP GETTING STUCK ON THIS SIMPLE PROBLEM< HELP: Simplify:

OpenStudy (anonymous):

give me the problem and i shall help

OpenStudy (iwanttogotostanford):

\[-2\times3-15\div(-5)4-(-7)\]

OpenStudy (iwanttogotostanford):

I keep ending up with: -21 divided by 140 but i Know thats not right

OpenStudy (anonymous):

13

OpenStudy (iwanttogotostanford):

i need to know the steps you took because I keep doing it over and get the wrong result

OpenStudy (anonymous):

Lol i have 2 tabs of openstudy open i typed 13 in the other one by mistake lols

OpenStudy (anonymous):

Want me to refer you to a lesson with good explanations Because im bad at explaining things im sorry :C

OpenStudy (anonymous):

(−2)(3)− 15 −5 (4)−(−7) =−6− 15 −5 (4)−(−7) =−6−(−3)(4)−(−7) =−6−(−12)−(−7) =6−(−7) =13

OpenStudy (anonymous):

thats my work

OpenStudy (iwanttogotostanford):

ok i usualyy am good at these but not this one

OpenStudy (anonymous):

Thats k wanna medal me

OpenStudy (iwanttogotostanford):

ok

OpenStudy (anonymous):

Thx if you got anymore question tag me in them

OpenStudy (anonymous):

To tag put a @symbol then my name

OpenStudy (anonymous):

@Lord_Frieza example

OpenStudy (iwanttogotostanford):

ok